# algorithmic

### v1

@@ 1 Introduction
The advent of deep generative models has revolutionized the field of image generation. Key developments marked by Variational Autoencoders (Kingma & Welling, 2013), Generative Adversarial Networks (Goodfellow et al., 2014), and the more emergent diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015), have demonstrated an unprecedented capability in producing high-quality images. They have been followed by remarkable results in applications such as super-resolution (Ledig et al., 2017; Li et al., 2022), face generation (Karras et al., 2019; 2020; 2021), and text-to-image generation (Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022). However, despite these advancements, the critical challenge of balancing fidelity (the realism of the generated images) and diversity (the variety in the generated images) remains unaddressed, despite this ability being essential for the practical applicability of these models. Our work focuses on tackling this crucial need for improved control mechanisms in generative models.
@@ 2 Background and Related Work
Figure 1: Images generated by StyleGAN2 (Top), ProjectedGAN (Middle), and IDDPM (Bottom). The training datasets are FFHQ, LSUN-Church, and LSUN-Bedroom, respectively. For each row, the left four images obtained the highest pseudo density out of 1000 generations, and the right four obtained the lowest.
@@ 3 Pseudo Density
Our control approach aims to explicitly manipulate the occurring likelihood of data samples, ideally based on their probability density in the real data distribution. However, estimating the density of high-dimensional data such as images is not straightforward, primarily due to the complexities introduced by the high dimensionality and the fact that image data exist in a low-dimensional manifold. In this section, we introduce an effective metric, termed pseudo density, for estimating the density of image data. Specifically, we first employ a pre-trained image feature extractor to lower the dimensionality of the real image data while capturing the semantics. Then, for each real sample, the average distance $d$ to its nearest neighbors is calculated. Based on the average neighbor distances, we can estimate the volume “occupied” by each sample, and the density is inversely proportional to the volume. Finally, we train a light-weight fully-connected network to fit the estimated density given the extracted feature as input. To compute the pseudo density of any generated sample, we feed it into the cascade of the feature extractor and the trained network. Note that this pseudo density metric differs from the Fréchet Inception Distance (FID) (Heusel et al., 2017): Pseudo density evaluates individual samples while the FID evaluates the overall performance of the generative model by calculating the distributional distance between the real and generated images. We provide more details on these steps in the following.
@@
Estimating the real data distribution. We estimate the real data distribution in a feature space, using the extracted features of samples from the dataset. First, the features of all real samples in the dataset can be extracted by a pre-trained feature extractor, such as a Vision Transformer (ViT) (Dosovitskiy et al., 2020) trained on ImageNet (Deng et al., 2009). Then, for each data sample $x_{i}$, we find its top-$k$ nearest neighbors based on the Euclidean distance in the feature space and compute its average distance to these $k$ samples $d_{i}^{k}$. With the assumption that the sample and its neighbors are locally uniformly distributed within a sphere of radius $d$, we can compute the volume a sample “occupies” as $V_{i}^{k}=C\cdot{(d_{i}^{k})}^{n}$, where $n$ is the dimensionality of the feature space and $C$ is a constant only conditioned on $n$. The density at the data point $x_{i}$ is hence inversely proportional to the volume, i.e., $\hat{\rho}_{i}^{k}\propto\frac{1}{V_{i}^{k}}$. Finally, for numerical stability, we compute the normalized density across all data points as $\rho_{i}=N\cdot\hat{\rho}_{i}^{k}\cdot\frac{1}{\sum^{N}_{i}\hat{\rho}_{i}^{k}}$, where $N$ is the total number of samples. Note that although image data (as well as their extracted features) are in a high-dimensional space, they reside in a manifold of much lower dimensionality. Hence, in practice, we set $n$ to be a small number (e.g., 1 or 2), preventing the volume $V$ from the curse of dimensionality.
@@
Learning a model that predicts the density. To efficiently compute the density for any data point $x$ and enable gradient back-propagation, we train a simple regression network $\rho(\cdot)$ to fit the density ${\rho_{0},\rho_{1},...,\rho_{N-1}}$ with the extracted features of real samples as input ${F(x_{0}),F(x_{1}},...,F(x_{N-1})$. We observed that a basic fully-connected network consisting of only three hidden layers with fewer than one million parameters is capable of fitting well to the real data while generalizing very effectively. We term the output of the network as pseudo density. Compared to directly computing the density of a generated sample by injecting it into all real samples, this approach has lower computational overhead and ensures that the gradient can be back-propagated from the metric to features.
@@ 4 Density-Based Perturbation
In the context of image synthesis, such a perturbation optimizes the latent vector within its neighborhood such that it leads to a higher-density or lower-density generated image, hence increasing the realism or uniqueness of the generated image without drastically changing the image content. Let us consider a randomly-sampled latent vector $\bm{z}$. We explore the neighborhood of $\bm{z}$ using a perturbation $\delta$ based on the pseudo density. To achieve this, we employ the PGD attack (Madry et al., 2018) strategy to minimize the pseudo density of the features of $G(\bm{z}+\bm{\delta})$ under the constraint $\|\bm{\delta}\|_{\infty}\leq\epsilon$, where $\epsilon$ is the predefined the adversarial budget. The pseudo-code of our per-sample perturbation method is provided in Algorithm 1. In Madry et al. (2018), the perturbation is applied directly to the image and the magnitude of $\epsilon$ should be kept small to ensure that the changes are imperceptible in the image; in our context, however, the perturbation is applied to the latent code and the resulted changes in the image are intended to be noticeable. Additionally, $\epsilon$ determines the magnitude of the pseudo density shift and the extent of the content changes in the generated image. A larger $\epsilon$ yields a large change in pseudo density and also more significant changes in the content and realism/uniqueness of the generated image.
@@ 5.1 Inference with Importance Sampling
Figure 3: Precision-recall trade-off. For all methods, the precision and recall metrics compete with each other. For polarity sampling, we used the values reported in Humayun et al. (2022a) due to replication difficulty. For our method, we adopted different density thresholds $\tau$ for each dashed line with varying importance weights $w$ ranging from $0.01$ to $100$. For better visualization, we only report part of the results for polarity sampling and truncation.
@@ 5.2 Fine-tuning with Importance Sampling
In contrast to the previous method that manipulates the generative data distribution during inference time, fine-tuning a generative model achieves it during training time. Compared to the previous method, fine-tuning does not introduce any computational overhead during inference. Furthermore, it enables the realism/uniqueness adjustment of any generated images, given their latent vectors. Similarly to the inference-time method, fine-tuning with importance sampling requires the same hyperparameters, the density threshold $\tau$ and the importance weight $w$, to yield different controlling effects. Before fine-tuning, we compute the pseudo density for every sample in the dataset. We can also determine the density threshold $\tau$ based on the densities of all samples. During our fine-tuning procedure, for each iteration, training examples are sampled from the dataset via the importance sampling strategy described before based on the pre-computed pseudo density.
In general, our fine-tuning approach does not modify the training algorithm except for how real samples are sampled from the dataset in each iteration. Therefore, it is compatible with any generative models including GANs and diffusion models. Moreover, it can be further improved for certain generative models such as GANs, in which we have not only the real data distribution $p_{r}(x)$ but also the generated data distribution $p_{g}(x)$ to play with during training. Specifically, when the generator and the discriminator reach their equilibrium, we have $\frac{p_{r}(x)}{p_{r}(x)+p_{g}(x)}=\frac{1}{2}$ (Goodfellow et al., 2014), and thus we obtain the generated data distribution $p_{g}(x)=p_{r}(x)$. In the case where we only perform importance sampling to the real data distribution $p_{r}(x)$, we have $p_{g}(x)=\textit{Pert}_{r}(p_{r}(x))$ upon convergence, where we use $\textit{Pert}_{g}(\cdot)$ to represent the perturbation function that maps the original distribution to the sampled one. When we also apply importance sampling to the generated data distribution, the equilibrium becomes $\textit{Pert}_{g}(p_{g}(x))=\textit{Pert}_{r}(p_{r}(x))$. By performing importance sampling on both $p_{g}(x)$ and $p_{r}(x)$ with different importance weights, we can achieve a stronger control over the generated data distribution $p_{g}(x)$. A practical choice is to set these weights as reciprocal pairs, as in our experiments. We provide the pseudo-code for the method in Algorithm 3.
@@
Additionally, our fine-tuning strategy is compatible with inference-time methods such as truncation and the proposed importance sampling strategy outlined in Section 5, enabling a combined enhancement over fidelity and diversity when applied in conjunction, as demonstrated in the Appendix A.4.
6 Experiments
In Figure 2, we showcase examples of images generated by the pre-trained models to illustrate the effects of our proposed per-sample perturbation strategy, which achieves large density change while preserving content of the images. By applying perturbation that increases density, the generated images exhibit greater realism, characterized by simplified backgrounds, fewer objects, etc. By contrast, perturbation to lower density leads to images with higher uniqueness, featuring more complex backgrounds, the presence of rare features in the datasets, etc. We compare the effectiveness of our method with that of a random perturbation (which adds a random vector to the latent vector) in the appendix section A.2, where additional examples are provided as well.
@@ 6 Experiments
In Figure 3, we demonstrate the trade-off between precision (fidelity) and recall (diversity) achieved by density-based importance sampling during inference. Our approach achieves a better overall precision-recall trade-off than the other two inference-time methods, polarity sampling (Humayun et al., 2022a) and truncation (Karras et al., 2019), especially in the improved-precision area. It’s important to note that polarity sampling may face compatibility issues with diffusion models due to the computational complexity of the denoising process. The truncation method is adaptable to the intermediate $\mathcal{W}+$ space for StyleGAN (Karras et al., 2019; 2020; 2021) models, but is generally limited to the input $\mathcal{Z}$ space for others, which could result in suboptimal performance. For instance, for Improved DDPM, the truncation method struggles around the pre-trained trade-off point. Unlike these methods, neither the computational complexity nor particular model architectures restricts the applicability of our method since our approach directly manipulate the output data.
@@ A.2 Discriminator Output versus Pseudo Density
Many works, including LOGAN (Wu et al., 2019), discriminator rejection sampling (Azadi et al., 2018), and discriminator optimal transport (Tanaka, 2019), have explored the output of a trained discriminator as a metric for sample realism. Ideally, discriminator output may indicate sample realism in the context of Generative Adversarial Networks since the discriminator learns to tell if a sample is real or generated. Azadi et al. (2018) demonstrated that discriminator output aligns well with realism in a 2D toy dataset. However, our experiments suggested that it works less ideally for image data, which are in much higher dimensions. In Figure 5, we showcased the results of per-sample perturbation based on pseudo density and discriminator output. Additionally, we compared to random perturbation, which replaces the gradient-descent direction with a random direction while keeping the same magnitude in each step of the PGD attack.
@@ A.3 Extra Computational Overhead
The proposed rejection sampling strategies, as described in Section 5, Algorithm 2 and Algorithm 3, result in extra computational cost since only a proportion of the generated samples are accepted. Given the importance weight $w<1$ and let the density threshold be $p\%$ of estimated densities of the data, assuming the generated data perfectly fit the real data, a generated sample is accepted with probability $0.01p+(1-0.01p)\times w$ on average, hence the computational cost is approximately $\frac{1}{0.01p+(1-0.01p)\times w}$ times as before. The case where $w>1$ can be handled similarly. In our experiments, generating samples via StyleGAN2 with $p=50,w=0.01$ slows down from the original $13$ sec/kimgs to $26$ sec/kimgs while generating with $p=20,w=0.01$ slows down to $60$ sec/kimgs. Fine-tuning StyleGAN2 with $p=80,w=33.0$ drops from $56$ sec/kimgs to $74$ sec/kimgs. However, fine-tuning diffusion models does not suffer from obvious slower training speed since it only involves importance sampling on the real data where rejection sampling is not applied.

### v2

@@ 1 Introduction
The advent of deep generative models has revolutionized the field of image generation. Key developments marked by Variational Autoencoders (Kingma & Welling, 2013), Generative Adversarial Networks (Goodfellow et al., 2014), and the more emergent diffusion models (Ho et al., 2020; Sohl-Dickstein et al., 2015), have demonstrated an unprecedented capability in producing high-quality images. They have been followed by remarkable results in applications such as classification (Xu et al., 2021; Xu & Le, 2022; Xu et al., 2023a; b), super-resolution (Ledig et al., 2017; Li et al., 2022), face generation (Karras et al., 2019; 2020; 2021), and text-to-image generation (Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022). However, despite these advancements, the critical challenge of balancing fidelity (the realism of the generated images (Xu et al., 2024)) and diversity (the variety in the generated images) remains unaddressed, despite this ability being essential for the practical applicability of these models. Our work focuses on tackling this crucial need for improved control mechanisms in generative models.
@@ 2 Background and Related Work
Figure 1: Overview of the proposed pseudo density metric. (a) The pseudo-density of each sample is inversely proportional to its average distance to its K-Nearest Neighbors in a feature space. (b) We use these pseudo density to train a density prediction network, which then can be used on any generated samples (c).
Figure 2: Images generated by StyleGAN2 (Top), ProjectedGAN (Middle), and IDDPM (Bottom). The training datasets are FFHQ, LSUN-Church, and LSUN-Bedroom, respectively. For each row, the left four images obtained the lowest pseudo density out of 1000 generations, and the right four obtained the highest.
@@ 3 Pseudo Density
Our control approach aims to explicitly manipulate the occurring likelihood of data samples, ideally based on their probability density in the real data distribution. However, estimating the density of high-dimensional data such as images is not straightforward, primarily due to the complexities introduced by the high dimensionality and the fact that image data exist in a low-dimensional manifold. In this section, we introduce an effective metric, termed pseudo density, for estimating the density of image data. Specifically, we first employ a pre-trained image feature extractor to lower the dimensionality of the real image data while capturing the semantics. Then, for each real sample, the average distance $d$ to its nearest neighbors is calculated. Based on the average neighbor distances, we can estimate the volume “occupied” by each sample, and the density is inversely proportional to the volume. Finally, we train a light-weight fully-connected network to fit the estimated density given the extracted feature as input. To compute the pseudo density of any generated sample, we feed it into the cascade of the feature extractor and the trained network. Note that this pseudo density metric differs from the Fréchet Inception Distance (FID) (Heusel et al., 2017): Pseudo density evaluates individual samples while the FID evaluates the overall performance of the generative model by calculating the distributional distance between the real and generated images. We illustrate the steps of computing pseudo density in Figure 1 and provide more details in the following.
@@
Estimating the real data distribution. We estimate the real data distribution in a feature space, using the extracted features of samples from the dataset. First, the features of all real samples in the dataset can be extracted by a pre-trained feature extractor, such as a Vision Transformer (ViT) (Dosovitskiy et al., 2020) trained on ImageNet (Deng et al., 2009). Then, for each data sample $x_{i}$, we find its top-$k$ nearest neighbors based on the Euclidean distance in the feature space and compute its average distance to these $k$ samples $d_{i}^{k}$. With the assumption that the sample and its neighbors are locally uniformly distributed within a sphere of radius $d$, we can compute the volume a sample “occupies” as $V_{i}^{k}=\frac{\pi^{n/2}}{\Gamma(\frac{n}{2}+1)}{(d_{i}^{k})}^{n}=C\cdot{(d_{i}^{k})}^{n}$, where $n$ is the dimensionality of the feature space, $\Gamma$ is Euler’s gamma function, and $C$ is a constant only conditioned on $n$. The density at the data point $x_{i}$ is hence inversely proportional to the volume, i.e., $\hat{\rho}_{i}^{k}\propto\frac{1}{V_{i}^{k}}$. Finally, for numerical stability, we compute the normalized density across all data points as $\rho_{i}=N\cdot\hat{\rho}_{i}^{k}\cdot\frac{1}{\sum^{N}_{i}\hat{\rho}_{i}^{k}}$, where $N$ is the total number of samples. Note that although image data (as well as their extracted features) are in a high-dimensional space, they reside in a manifold of much lower dimensionality. Hence, in practice, we set $n$ to be a small number (e.g., 1 or 2), preventing the volume $V$ from the curse of dimensionality.
@@
Learning a model that predicts the density. To efficiently compute the density for any data point $x$ and enable gradient back-propagation, we train a simple regression network $\rho(\cdot)$ to fit the density ${\rho_{0},\rho_{1},...,\rho_{N-1}}$ with the extracted features of real samples as input ${F(x_{0}),F(x_{1})},...,F(x_{N-1})$. We observed that a basic fully-connected network consisting of only three hidden layers with fewer than one million parameters is capable of fitting well to the real data while generalizing very effectively. We term the output of the network as pseudo density. Compared to directly computing the density of a generated sample by injecting it into all real samples, this approach has lower computational overhead and ensures that the gradient can be back-propagated from the metric to features.
@@ 4 Density-Based Perturbation
In the context of image synthesis, such a perturbation optimizes the latent vector within its neighborhood such that it leads to a higher-density or lower-density generated image, hence increasing the realism or uniqueness of the generated image without drastically changing the image content. Let us consider a randomly-sampled latent vector $\bm{z}$. We explore the neighborhood of $\bm{z}$ using a perturbation $\delta$ based on the pseudo density. To achieve this, we employ the PGD attack (Madry et al., 2018) strategy to minimize the pseudo density of the features of $G(\bm{z}+\bm{\delta})$ under the constraint $\|\bm{\delta}\|_{\infty}\leq\epsilon$, where $\epsilon$ is the predefined adversarial budget. The pseudo-code of our per-sample perturbation method is provided in Algorithm 1. In Madry et al. (2018), the perturbation is applied directly to the image and the magnitude of $\epsilon$ should be kept small to ensure that the changes are imperceptible in the image; in our context, however, the perturbation is applied to the latent code and the resulted changes in the image are intended to be noticeable. Additionally, $\epsilon$ determines the magnitude of the pseudo density shift and the extent of the content changes in the generated image. A larger $\epsilon$ yields a large change in pseudo density and also more significant changes in the content and realism/uniqueness of the generated image.
@@ 5.1 Inference with Importance Sampling
Figure 4: Precision-recall trade-off. For all methods, the precision and recall metrics compete with each other. For polarity sampling, we used the values reported in Humayun et al. (2022a) due to replication difficulty. For our method, we adopted different density thresholds $\tau$ for each dashed line with varying importance weights $w$ ranging from $0.01$ to $100$. For better visualization, we only report part of the results for polarity sampling and truncation. We visualize the Pareto frontier for polarity sampling and ours, illustrating the optimal trade-offs achieved across various hyperparameter configurations.
@@ 5.2 Fine-tuning with Importance Sampling
In contrast to our previously introduced method, which manipulates the generated data distribution at inference time, the current method employs importance sampling during the fine-tuning process to adjust the data distribution at training time. As a result, fine-tuning does not introduce any computational overhead during inference. Furthermore, it enables the realism/uniqueness adjustment of any generated images, given their latent vectors. Similarly to the inference-time method, fine-tuning with importance sampling requires the same hyperparameters, the density threshold $\tau$ and the importance weight $w$, to yield different controlling effects. Before fine-tuning, we compute the pseudo density for every sample in the dataset, and then determine the density threshold $\tau$ based on the computed densities of all samples. During fine-tuning, training examples are sampled from the dataset using the importance sampling strategy described before based on the pre-computed pseudo density.
@@
In general, our fine-tuning approach does not modify the training algorithm except for how images are sampled from the data distribution in each iteration. Therefore, it is compatible with any generative models including GANs and diffusion models. Moreover, it can be further enhanced for certain generative models, particularly GANs, for which importance sampling can be applied not only to the real data distribution $p_{r}(x)$, but also to the generated data distribution $p_{g}(x)$, which serves as input to the discriminator during training. We provide the pseudo-code for the method for GANs in Algorithm 3. Specifically, when the generator and the discriminator reach their equilibrium, we have $\frac{p_{r}(x)}{p_{r}(x)+p_{g}(x)}=\frac{1}{2}$ (Goodfellow et al., 2014), and thus the learned generated data distribution $p_{g}(x)=p_{r}(x)$. In the case where we perform importance sampling to the real data distribution $p_{r}(x)$, we have $p_{g}(x)=\textit{Pert}_{r}(p_{r}(x))$ upon convergence, where we use $\textit{Pert}_{r}(\cdot)$ to represent the perturbation function that maps the original distribution to the sampled one. If we also apply importance sampling to the generated data distribution during training, the equilibrium becomes $\textit{Pert}_{g}(p_{g}(x))=\textit{Pert}_{r}(p_{r}(x))$. This allows a stronger control over the learned generated data distribution $p_{g}(x)$ after training. A practical choice is to perform importance sampling on $p_{g}(x)$ and $p_{r}(x)$ with reciprocal importance weights $w$ and $\frac{1}{w}$, as in Algorithm 3.
@@
Additionally, our fine-tuning strategy is compatible with inference-time techniques, including truncation and the proposed importance sampling strategy detailed in Section 5. When inference-time methods are applied in conjunction with our fine-tuning strategy, a stronger enhancement over fidelity or diversity can be achieved. Empirical evidence supporting this combined effect is presented in Appendix A.4.
@@
6 Experiments
In Figure 3, we showcase examples of images generated by the pre-trained models to illustrate the effects of our proposed per-sample perturbation strategy, which achieves large density change while preserving content of the images. By applying perturbation that increases density, the generated images exhibit greater realism, characterized by simplified backgrounds, fewer objects, etc. By contrast, perturbation to lower density leads to images with higher uniqueness, featuring more complex backgrounds, the presence of rare features in the datasets, etc. In Appendix  A.2, we consider perturbation that increases the output of a GAN’s discriminator and demonstrate its inferior performance compared to pseudo density. Additionally, we compare our approach with Discriminator Rejection Sampling (Azadi et al., 2018) and Discriminator Optimal Transport (Tanaka, 2019), both of which rely on the trained discriminator’s output in GANs as an indicator for image realism.
@@ 6 Experiments
In Figure 4, we demonstrate the trade-off between precision (fidelity) and recall (diversity) achieved by density-based importance sampling during inference. Our approach achieves a better overall precision-recall trade-off than the other two inference-time methods, polarity sampling (Humayun et al., 2022a) and truncation (Karras et al., 2019), especially in the improved-precision area. It’s important to note that polarity sampling may face compatibility issues with diffusion models due to the computational complexity of the denoising process. The truncation method is adaptable to the intermediate $\mathcal{W}+$ space for StyleGAN (Karras et al., 2019; 2020; 2021) models, but is generally limited to the input $\mathcal{Z}$ space for others, which could result in suboptimal performance. For instance, for Improved DDPM, the truncation method struggles around the pre-trained trade-off point. Unlike these methods, neither the computational complexity nor particular model architectures restricts the applicability of our method since our approach directly manipulates the output data.
@@ A.2 Discriminator Output versus Pseudo Density
Many works, including LOGAN (Wu et al., 2019), Discriminator Rejection Sampling (Azadi et al., 2018), and Discriminator Optimal Transport (Tanaka, 2019), have explored the output of a trained discriminator as a metric for sample realism. Ideally, discriminator output may indicate sample realism in the context of Generative Adversarial Networks since the discriminator learns to tell if a sample is real or generated. DRS (Azadi et al., 2018) and DOT (Tanaka, 2019) demonstrated that discriminator output aligns well with realism in low-dimensional datasets and their methods improve FID for SNGAN (Miyato et al., 2018) and SAGAN (Zhang et al., 2019). However, our experiments suggested that it works badly for higher-resolution image data and larger modern GAN architectures such as StyleGAN (Karras et al., 2019; 2020). In Figure 8, we showcased the results of per-sample perturbation based on pseudo density and discriminator output using StyleGAN2 on FFHQ dataset. Additionally, we compared to random perturbation, which replaces the gradient-descent direction with a random direction while keeping the same magnitude in each step of the PGD attack.
We implemented DRS and DOT for StyleGAN2 and followed the original hyperparameters reported in the papers. As shown in Figure 7, where we plot the precision and recall for DRS and DOT alongside our approach and others, both DRS and DOT achieved performances similar to using the pre-trained model naively. Additionally, DRS and DOT obtained FID scores of 2.95 and 3.44, respectively, which are worse than that of the pretrained. These results indicate that discriminator-based methods like DRS and DOT do not achieve ideal precision-recall trade-off for high-resolution image data and larger, modern GAN models.
Figure 6: Precision-recall trade-off for StyleGAN2 on FFHQ.
Figure 7: Precision-recall trade-off for VQGAN on FFHQ.
@@ A.3 Extra Computational Overhead
The proposed rejection sampling strategies, as described in Section 5, Algorithm 2 and Algorithm 3, result in extra computational cost since only a proportion of the generated samples are accepted. Given the importance weight $w<1$ and let the density threshold be $p\%$ of estimated densities of the data, assuming the generated data perfectly fit the real data, a generated sample is accepted with probability $0.01p+(1-0.01p)\times w$ on average, hence the computational cost is approximately $\frac{1}{0.01p+(1-0.01p)\times w}$ times as before. The case where $w>1$ can be handled similarly.
In our experiments, generating samples using StyleGAN2 with importance sampling ($p=50,w=0.01$) slowed down from the original $13$ sec/kimgs to $26$ sec/kimgs while generating with $p=20,w=0.01$ slowed down to $60$ sec/kimgs. In contrast, generating samples with DRS (Azadi et al., 2018) and DOT (Tanaka, 2019) slowed down to $23$ sec/kimgs and $290$ sec/kimgs respectively. We were unable to obtain statistics for polarity sampling (Humayun et al., 2022a) due to difficulties in replicating the method, but the authors reported that preprocessing for StyleGAN2 took a daunting 14 days. In comparison, our method and DOT took several minutes, while DRS took less than an hour.
On the other hand, our fine-tuning method maintains the original inference efficiency but compromises training efficiency. Fine-tuning StyleGAN2 using importance sampling with $p=80,w=33.0$ lowered the speed from $56$ sec/kimgs to $74$ sec/kimgs. However, fine-tuning diffusion models did not exhibit a significant slowdown in training speed since it involves importance sampling only on the real data, avoiding the time-consuming rejection sampling. Notably, our fine-tuning process typically requires less than one hour to complete due to the small number of iterations needed.